# Quantum framework for Reinforcement Learning: integrating Markov Decision Process, quantum arithmetic, and trajectory search

### v1

@@ IV.3 Quantum trajectories search
We formulated the interaction between a quantum agent and its environment over $H$ time steps (Eq. (12)). A quantum trajectory here refers to the series of quantum states and actions the agent takes, together with the environment’s responses. Each trajectory |$\ket{T}$ can be represented as |$\ket{T}=\ket{s_{0},a_{0},s^{\prime}_{0},r_{0},s_{1},a_{1},s^{\prime}_{1},r_{1},\ldots,s_{H-1},a_{H-1},s^{\prime}_{H-1},r_{H-1},g}$ where $s$ is the current state at each time step from $t=0$ to $t=H-1$, $a$ is the action taken by the agent, $s^{\prime}$ is the next state reached after taking an action at state $s$, $r$ is the reward received and $g$ is the total return throughout the trajectory.
@@
V Experiments and Results
Figure 3: Graphical representation of a classical Markov Decision Process (MDP) with four states $(s_{0},s_{1},s_{2},s_{3})$, two actions $(a_{0},a_{1})$ and rewards $(r_{0},r_{1},r_{2},r_{3})$. The arrows between states indicate the transitions associated with each action. Transition probabilities are labeled on each arrow.
In the experiments and results section of this work, we first present the dynamics of the classical Markov Decision Process (MDP) through a graphical representation of state transitions and associated probabilities, as shown in Fig. 3. The diagram includes four states labeled $s_{0},s_{1},s_{2}$ and $s_{3}$, and two actions, $a_{0}$ and $a_{1}$. The arrows connecting the states indicate the transitions between these states for a given action. Each transition is labeled with a probability value that denotes the likelihood of moving from one state to another under a specific action. These probabilities provide insights into the stochastic nature of the state transitions. For example, transitioning from $s_{0}$ to $s_{1}$ with action $a_{0}$ has a probability of 0.6. Similarly, transitioning from $s_{0}$ to $s_{2}$ under action $a_{0}$ occurs with a probability of 0.4. The diagram also illustrates several self-transitions, such as the one in state $s_{3}$, where action $a_{1}$ leads to a transition back to $s_{3}$ with a probability of 1. These self-transitions capture situations where the system remains in the same state regardless of the action taken. The figure visualizes the probabilistic behavior of the system under different actions, as well as how the state evolves over time based on these actions.
This classical MDP serves as a reference framework, laying the groundwork for comparing it to the quantum version in our proposed QMDP model. The quantum model introduces superposition effects, enhancing the classical decision-making process with additional computational capabilities.
V.1 Experiment on quantum implementation of MDP
@@ V.1.1 Agent and environment on quantum computer
The circuit begins by preparing the superpositions of states and actions on quantum registers, representing the starting conditions of the agent-environment interaction. Rotation gates $R_{y}(\theta)$ represent the state transition probabilities, analogous to how probabilities govern transitions in the classical MDP. These gates parameterize the probability of transitioning between states, but in the quantum domain, they introduce quantum amplitudes instead of classical probabilities, allowing the agent to explore multiple potential states at once, due to quantum superposition. The controlled part of the rotation ensures that state transitions occur conditionally. The control qubits, representing the current state and action, dictate whether the target qubits (the next state registers) undergo the rotation $R_{y}(\theta)$. This conditional operation ensures that state transition probabilities are applied selectively, based on the current state-action pairs. Mathematically, the rotation angle $\theta$ is determined by the transition probability $P(s^{\prime}|s,a)$, encoding it into the quantum amplitudes of the next state registers. CNOT gates model the reward mechanism. Specifically, the CNOT gates connect the next state registers to the reward qubits, ensuring that the reward is applied conditionally, depending on the resulting state.
By implementing the classical MDP into the quantum domain, the quantum model allows the agent to explore and evaluate multiple state-action pairs, as well as their potential next states and rewards simultaneously. This exploration enables the agent to capture complex environment dynamics and offers potential improvements in computational efficiency. The transformations in this quantum model align with the principles of quantum mechanics to enhance the exploration of the state space compared to classical methods.
@@ V.1.2 Agent and environment single interaction on quantum computer
V.2 Experiment on agent-environment quantum interactions over multiple time steps
@@
Now, we present the classical measurement outcomes obtained from simulating the agent-environment interactions over 3 time steps. The classical measurement phase provides insight into the specific quantum trajectories generated by the QMDP circuit, reflecting the different paths the agent took based on its actions and the corresponding state transitions. In this experiment, we used the IBM quantum simulator to perform this measurement. Figure 9 shows the quantum trajectory distribution over 3 time steps. Table 4 (see Appendix A) presents the list of quantum trajectories along with their corresponding trajectory numbers. Each quantum trajectory is represented as a binary string, which corresponds to the sequences of state-action-next state-reward and the return which are observed during the simulation of the interactions over 3 time steps. The y-axis displays the total count of occurrences for each trajectory. Trajectories that occur more frequently are represented by higher bars, showing which paths the agent most followed during these interactions.
@@
V.3 Experiment on quantum trajectories search
In this section, we present the results of implementing Grover’s search algorithm to identify optimal quantum trajectories within the quantum reinforcement learning (QRL) framework outlined in the method section. The objective of this experiment is to efficiently search through the set of possible quantum trajectories generated from agent-environment interactions to find those that maximize the overall return, thereby guiding the discovery of optimal policies for the QRL agent. Each trajectory is composed of a sequence of quantum states, actions, next state, rewards, and a return value, and Grover’s search algorithm is employed to efficiently identify those trajectories with the highest return from a large set of possible agent-environment interactions.
@@ V.3 Experiment on quantum trajectories search
The following provides a detailed analysis of the results for first scenario, illustrating the performance of Grover’s search algorithm in identifying the optimal quantum trajectories that yield the maximum return. These results demonstrate the potential of quantum algorithms to enhance trajectory search processes in reinforcement learning tasks. This experiment not only demonstrates the application of Grover’s search algorithm but also highlights the significant advantages gained from implementing a classical Markov decision process (MDP) into the quantum domain. By manipulating the quantum version of MDP, we can efficiently search through possible trajectories and state-action pairs using Grover’s search in only one call to the oracle. This quantum-enhanced framework provides a substantial computational advantage over classical methods, which would typically require more computational resources to explore the same search space.
@@
In the first case, we examine the scenario where the agent starts from a fixed initial state $s_{0}$ and terminates at state $s_{3}$, aiming to achieve the maximum return over three time steps. The result of this experiment is visualized in Fig. 10, which shows the distribution of quantum trajectories identified by Grover’s search algorithm, as well as their respective counts based on how frequently each trajectory was sampled. The numbers along the x-axis represent distinct quantum trajectories numbers, each representing a unique sequences of states, actions, next states, rewards and return over 3 time steps (see Table 4 in Appendix A for the corresponding quantum trajectories). The bar illustrates the total count of each trajectory sampled during the measurement process after executing Grover’s search.
@@ VI Discussion
The experimental results validate that the quantum implementation of the Markov Decision Process (QMDP) not only aligns with classical MDP dynamics but also effectively leverages quantum properties to enhance performance. Our quantum model maintains the core structure of state transitions and action choices as observed in the classical MDP and ensures that the decision-making of QMDP remain consistent with that of classical MDP. This validates that our quantum approach can successfully replicate the behavior of classical MDP while introducing potential improvements in speed and computational efficiency. Another key observation is that with only 3 time steps, the quantum agent can effectively explore the possible trajectories and observe the optimal policies for each state. The quantum agent’s ability for simultaneous exploration of multiple trajectories, resulting in a faster and more efficient determination of the optimal policy compared to traditional algorithms. Furthermore, the use of Grover’s search algorithm within our quantum MDP framework demonstrates that a single call to the oracle is sufficient to identify the desired trajectories. Grover’s algorithm offers a speedup, allowing the quantum agent to perform optimal trajectory searches with substantially fewer computational steps. This efficient solution highlights an improvement in speed and computational resources compared to classical search algorithms. These advantages highlight the potential of quantum computing in complex search and optimization processes, providing faster solutions with reduced computational resources.
@@ Acknowledgements.
This work was supported by JST COI-NEXT Grant Number JPMJPF2221 and the Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@ Appendix B Results for Quantum Trajectories Search
In this section, we examine Grover’s search for identifying the optimal trajectories for broader search space where the agent can start from any initial state within the state space with equal probability and must terminate at state $s_{3}$, aiming to maximize the total return over 3 time steps. This scenario expands the search space by introducing uncertainty in the agent’s starting state, making the task of finding the optimal trajectory more complex. The result of this experiment, visualized in Fig. 13, shows the distribution of quantum trajectories identified by Grover’s search algorithm. The trajectory numbers on the x-axis represent distinct quantum trajectories, with each number representing a unique sequences of state-action-next state and reward over 3 time steps, followed by a return value (see Table 4 in Appendix A for the corresponding quantum trajectories).

### v2

@@ IV.3 Quantum trajectories search
We formulated the interaction between a quantum agent and its environment over $T$ time steps (Eq. (12)). A quantum trajectory here refers to the series of quantum states and actions the agent takes, together with the environment’s responses. Each trajectory |$\ket{T}$ can be represented as |$\ket{T}=\ket{s_{0},a_{0},s^{\prime}_{0},r_{0},s_{1},a_{1},s^{\prime}_{1},r_{1},\ldots,s_{T-1},a_{T-1},s^{\prime}_{T-1},r_{T-1},g}$ where $s$ is the current state at each time step from $t=0$ to $t=T-1$, $a$ is the action taken by the agent, $s^{\prime}$ is the next state reached after taking an action at state $s$, $r$ is the reward received and $g$ is the total return throughout the trajectory.
@@
V Demonstrations and Results
In the demonstrations and results section of this work, we first present the dynamics of the classical Markov Decision Process (MDP) through a graphical representation of state transitions and associated probabilities, as shown in Fig. 3. The diagram includes four states labeled $s_{0},s_{1},s_{2}$ and $s_{3}$, and two actions, $a_{0}$ and $a_{1}$. The arrows connecting the states indicate the transitions between these states for a given action. Each transition is labeled with a probability value that denotes the likelihood of moving from one state to another under a specific action. These probabilities provide insights into the stochastic nature of the state transitions. For example, transitioning from $s_{0}$ to $s_{1}$ with action $a_{0}$ has a probability of 0.6. Similarly, transitioning from $s_{0}$ to $s_{2}$ under action $a_{0}$ occurs with a probability of 0.4. The diagram also illustrates several self-transitions, such as the one in state $s_{3}$, where action $a_{1}$ leads to a transition back to $s_{3}$ with a probability of 1. These self-transitions capture situations where the system remains in the same state regardless of the action taken. The figure visualizes the probabilistic behavior of the system under different actions, as well as how the state evolves over time based on these actions.
Figure 3: Graphical representation of a classical MDP with four states $(s_{0},s_{1},s_{2},s_{3})$, two actions $(a_{0},a_{1})$ and rewards $(r_{0},r_{1},r_{2},r_{3})$. The arrows between states indicate the transitions associated with each action. Transition probabilities are labeled on each arrow.
This classical MDP serves as a reference framework, laying the groundwork for comparing it to the quantum version in our proposed QMDP model. The quantum model introduces superposition effects, enhancing the decision-making process with additional computational capabilities. Specifically, the superposition principle allows the quantum model to represent and process all possible state-action combinations in parallel. This enables the agent to explore multiple trajectories simultaneously, significantly reducing the number of interactions required to learn optimal policies. The quantum implementation further benefits from Grover’s search algorithm, which provides a speedup for identifying optimal trajectories based on cumulative returns. These computational advantages go beyond classical capabilities by allowing efficient encoding, evaluation, and discovery of optimal sequences within the quantum domain.
V.1 Demonstration of quantum implementation of MDP
@@ V.1.1 Agent and environment on quantum computer
The circuit begins by preparing the superpositions of states and actions on quantum registers, representing the starting conditions of the agent-environment interaction. Rotation gates $R_{y}(\theta)$ represent the state transition probabilities, analogous to how probabilities govern transitions in the classical MDP. These gates parameterize the probability of transitioning between states, but in the quantum domain, they introduce quantum amplitudes instead of classical probabilities, allowing the agent to explore multiple potential states at once, due to quantum superposition. The controlled $R_{y}(\theta)$ rotations apply state transitions conditionally, such that only specific state-action pairs (encoded in the control qubits) activate the rotation on the next state registers. Mathematically, the rotation angle $\theta$ is determined by the transition probability $P(s^{\prime}|s,a)$, encoding it into the quantum amplitudes of the next state registers. CNOT gates model the reward mechanism. Specifically, the CNOT gates connect the next state registers to the reward qubits, ensuring that the reward is applied conditionally, depending on the resulting state.
By implementing the classical MDP in the quantum domain, the quantum model allows the agent to explore and evaluate multiple state-action pairs, as well as their potential next states and rewards, simultaneously. By leveraging quantum superposition, this exploration enables the agent to capture the complex dynamics of the environment, including its stochastic transitions and multiple possible outcomes in parallel, which would otherwise require repeated sampling in classical settings. As a result, the quantum model reduces the number of required interactions, leading to improved sample efficiency and potentially enhancing computational performance. Therefore, these transformations enhance the agent’s ability to explore the state space more effectively than classical methods.
@@ V.1.2 Agent and environment single interaction on quantum computer
V.2 Demonstration of agent-environment quantum interactions over multiple time steps
@@
Now, we present the classical measurement outcomes obtained from simulating the agent-environment interactions over 3 time steps. The classical measurement phase provides insight into the specific quantum trajectories generated by the QMDP circuit, reflecting the different paths the agent took based on its actions and the corresponding state transitions. In this demonstration, we used IBM Qiskit’s Aer simulator to perform this measurement (see Appendix C for detailed specifications). Figure 9 shows the quantum trajectory distribution over 3 time steps. Table 4 (see Appendix A) presents the list of quantum trajectories along with their corresponding trajectory numbers. Each quantum trajectory is represented as a binary string, which corresponds to the sequences of state-action-next state-reward and the return which are observed during the simulation of the interactions over 3 time steps. The y-axis displays the total count of occurrences for each trajectory. Trajectories that occur more frequently are represented by higher bars, showing which paths the agent most followed during these interactions.
@@
V.3 Demonstration of quantum trajectories search
In this section, we present the results of implementing Grover’s search algorithm to identify optimal quantum trajectories within the quantum reinforcement learning (QRL) framework outlined in the method section. The objective of this demonstration is to efficiently search through the set of possible quantum trajectories generated from agent-environment interactions to find those that maximize the overall return, thereby guiding the discovery of optimal policies for the QRL agent. Each trajectory is composed of a sequence of quantum states, actions, next state, rewards, and a return value, and Grover’s search algorithm is employed to efficiently identify those trajectories with the highest return from a large set of possible agent-environment interactions.
@@ V.3 Demonstration of quantum trajectories search
The following provides a detailed analysis of the results for first scenario, illustrating the performance of Grover’s search algorithm in identifying the optimal quantum trajectories that yield the maximum return. These results reveal the potential of quantum algorithms to enhance trajectory search processes in reinforcement learning tasks. This demonstration not only illustrates the application of Grover’s search algorithm but also highlights the significant advantages gained from implementing a classical Markov decision process (MDP) into the quantum domain. By manipulating the quantum version of MDP, we can efficiently search through possible trajectories and state-action pairs using Grover’s search in only one call to the oracle. This quantum-enhanced framework provides a substantial computational advantage over classical methods, which would typically require more computational resources to explore the same search space.
@@
In the first case, we examine the scenario where the agent starts from a fixed initial state $s_{0}$ and terminates at state $s_{3}$, aiming to achieve the maximum return over three time steps. The result of this demonstration is visualized in Fig. 10, which shows the distribution of quantum trajectories identified by Grover’s search algorithm, as well as their respective counts based on how frequently each trajectory was sampled. The numbers along the x-axis represent distinct quantum trajectories numbers, each representing a unique sequences of states, actions, next states, rewards and return over 3 time steps (see Table 4 in Appendix A for the corresponding quantum trajectories). The bar illustrates the total count of each trajectory sampled during the measurement process after executing Grover’s search.
@@
V.4 Advantages of quantum RL over classical approaches
A core benefit of our work lies in its ability to execute the entire reinforcement learning process within the quantum domain. The integration of quantum information formalism into reinforcement learning presents several significant advantages over classical approaches.
By representing both the agent and environment as quantum components and leveraging quantum superposition, the quantum model enables the simultaneous exploration of multiple trajectories. This parallelism allows the quantum agent to evaluate numerous interaction sequences in a single execution, significantly reducing the number of interactions required to explore the environment. As a result, the quantum approach achieves much greater sample efficiency compared to classical methods, which typically process one interaction sequence at a time and require many iterations to converge on an optimal policy.
Moreover, the implementation of Grover’s search algorithm enables efficient trajectory search with a single oracle call. This quantum-enhanced approach significantly accelerates the identification of high-return trajectories while avoiding the iterative evaluations required by classical methods. In our demonstration, the trajectories identified by the quantum approach match those found by classical Q-learning but are achieved with substantially lower computational cost. This highlights the effectiveness of Grover’s algorithm in delivering significant speedups in trajectory search compared to classical reinforcement learning algorithms. By integrating quantum superposition and Grover’s search, the proposed framework achieves both improved sample efficiency and substantial computational advantages over classical reinforcement learning.
@@ VI Discussion
The results validate that the quantum implementation of the Markov Decision Process (QMDP) not only aligns with classical MDP dynamics but also effectively leverages quantum properties to enhance performance. Our quantum model maintains the core structure of state transitions and action choices as observed in the classical MDP and ensures that the decision-making of QMDP remain consistent with that of classical MDP. This validates that our quantum approach can successfully replicate the behavior of classical MDP while introducing potential improvements in speed and computational efficiency. Another key observation is that with only 3 time steps, the quantum agent can effectively explore the possible trajectories and observe the optimal policies for each state. The quantum agent’s ability for simultaneous exploration of multiple trajectories, resulting in a faster and more efficient determination of the optimal policy compared to traditional algorithms. Furthermore, the use of Grover’s search algorithm within our quantum MDP framework demonstrates that a single call to the oracle is sufficient to identify the desired trajectories. Grover’s algorithm offers a speedup, allowing the quantum agent to perform optimal trajectory searches with substantially fewer computational steps. This efficient solution highlights an improvement in speed and computational resources compared to classical search algorithms. These advantages highlight the potential of quantum computing in complex search and optimization processes, providing faster solutions with reduced computational resources.
@@ VII Conclusion
This work serves as a foundational step in advancing the field of quantum reinforcement learning (QRL). It provides a concrete basis for future research aimed at scaling quantum-native RL systems, exploring more complex environments, and integrating with evolving quantum hardware platforms. Beyond its theoretical contributions, the proposed framework demonstrates strong potential for real-world applications. In autonomous driving, quantum MDP enables simultaneous evaluation of multiple driving trajectories, while Grover’s search accelerates optimal policy discovery, leading to faster decision-making for collision avoidance and route planning. In personalized healthcare, patient conditions and treatments can be modeled as MDPs, and the QRL framework evaluates multiple treatment plans in parallel, with Grover’s search identifying those with the highest reward for faster selection of effective therapies. In financial portfolio management, our approach enables parallel exploration of investment strategies, with Grover’s search rapidly identifying high-return paths, offering a significant computational advantage for real-time trading environments. These real-world scenarios illustrate how the proposed framework provides practical applicability and computational advantages, establishing a foundation for future advances in quantum reinforcement learning applications.
@@ Acknowledgements.
This work was supported by JST COI-NEXT Grant Number JPMJPF2221, JSPS KAKENHI Grant Number JP24K20843, and the Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@ Appendix B Results for Quantum Trajectories Search
In this section, we examine Grover’s search for identifying the optimal trajectories for broader search space where the agent can start from any initial state within the state space with equal probability and must terminate at state $s_{3}$, aiming to maximize the total return over 3 time steps. This scenario expands the search space by introducing uncertainty in the agent’s starting state, making the task of finding the optimal trajectory more complex. The result of this demonstration, visualized in Fig. 13, shows the distribution of quantum trajectories identified by Grover’s search algorithm. The trajectory numbers on the x-axis represent distinct quantum trajectories, with each number representing a unique sequences of state-action-next state and reward over 3 time steps, followed by a return value (see Table 4 in Appendix A for the corresponding quantum trajectories).
@@
Appendix C System and Simulation Environment Specifications
The computations in this work were conducted using the IBM Qiskit framework to simulate quantum circuits on a classical device. Below, we provide detailed specifications of both the quantum simulation environment and the classical hardware used in this study.
Quantum simulation environment specifications:
  1. (1)
    Quantum framework: IBM Qiskit (version 1.1.0)
  2. (2)
    Quantum simulator: IBM Qiskit Aer Simulator (version 0.14.2)
  3. (3)
    Simulation method: Statevector simulation
  4. (4)
    Simulation device: CPU-based simulation
  5. (5)
    Noise model: None (ideal simulation)
Classical device specifications:
  1. (1)
    Processor: AMD Ryzen 5 PRO 5650U with Radeon Graphics, 2.30 GHz
  2. (2)
    RAM: 16 GB
  3. (3)
    System type: 64-bit operating system, x64-based processor